OpenStudy (anonymous):

Prove the following statement: “If \[b[ \sqrt[3]{p}] + c [\sqrt[3]{p^2}]=0\] where a, b, c, p are rational and p is not a perfect cube, then a, b, c are all zero.”

OpenStudy (amistre64):

i dont see an "a"; but then again, proofs are not my cup o tea

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

Prove the following statement: “If \[a+b[p\sqrt[3]{3}]+c[\sqrt[3]{p}]=0\] where a, b, c, p are rational and p is not a perfect cube, then a, b, c are all zero.”

OpenStudy (anonymous):

Sorry again. I must sleep well. Prove the following statement: If \[a+b\sqrt[3]{p}+c\sqrt[3]{p^2}=0\] where a, b, c, p are rational and p is not a perfect cube, then a, b, c are all zero.”